OpenStudy (anonymous):

Solve. Find exact solutions: √(6+2x)-2=2

OpenStudy (anonymous):

start by adding \(2\) to both sides

OpenStudy (anonymous):

you get \(\sqrt{6+2x}=4\) square both sides and get \[6+2x=16\] then solve that one

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

what about this one\[\sqrt{4x-11}+2=\sqrt{6x-5}\] its really giving me trouble gecause its different. sorry for bothering